OpenStudy (anonymous):

please help (system of equations) Joanna went school supply shopping. She spent $30.96 on notebooks and pencils. Notebooks cost $2.37 each and pencils cost $1.20 each. She bought a total of 18 notebooks and pencils. How many of each did she buy? 6 notebooks; 12 pencils 13 notebooks; 5 pencils 8 notebooks; 10 pencils 11 notebooks; 7 pencils

OpenStudy (some.random.cool.kid):

to figure this out theres two ways ill give you the method i know, so she spent a total of 30.96 with 18 notebooks and pencils. this means means that you can take the answer and the prices for each item to figure out how many notebooks and pencils she purchased. Or you can do it another way where you try to put as many prices in until you get 30.96 which is the long and hard way.

OpenStudy (cwrw238):

you create 2 equations in x an y let x = number of notebooks and y the number of pencils there are a total of 18 so x + y = 18.......................(1) and the costs :- 2.37x + 1.2y = 30.96........(2) solve equations x(1) and (2) for x and y

OpenStudy (cwrw238):

* equations (1) and (2)

OpenStudy (some.random.cool.kid):

or instead you can use his hard method :P no offence.

OpenStudy (cwrw238):

its not so hard first multiply (1) by -1.2 and add the equations this will quickly give you the value of x

OpenStudy (some.random.cool.kid):

i know i was joking.

OpenStudy (cwrw238):

something seems to have frightened crayolaaa off, though....